Show step by step:

Solve for the inverse y= square root of x-2+4

1
Expert's answer
2021-03-30T13:44:44-0400

Given the function


"y =\sqrt{x-2}+4\\"


Let


"y=f(x)"

"\therefore f(x) =\sqrt{x-2}+4\\"

We want to find "f^{-1}(x)"

"y =\sqrt{x-2}+4\\
y-4 =\sqrt{x-2}\\"

Squaring both sides


"(y-4)^2 =(\sqrt{x-2})^2\\
(y-4)^2 = x-2\\
(y-4)^2+2 = x"

Therefore, the inverse of the function is


"f^{-1}(x) = (x-4)^2+2"